# Neural Signals Generate Clinical Notes in the Wild

### 06 Mar 2026 v2

@@ Appendix C Implementation Details
To facilitate reproducibility, we provide the complete source code for constructing the EEG–report benchmark from the Harvard Electroencephalography Database and the EHR database hosted on the Brain Data Science Platform, together with our model implementation, training scripts, data loaders, and pretrained weights at https://anonymous.4open.science/r/CELM-3AF4. Additionally, we provide the detailed implementation details of our approach in this section and the prompts used for the baselines in Appendix B. The dataset utilized for this study is publicly available and can be accessed at https://bdsp.io/content/harvard-eeg-db/4.1/ after obtaining credentialed access.
@@ C.1 Model Details and Hyperparameters
We use CBraMod (Wang et al., 2024b) as the EEG encoder and Qwen-4B-Instruct-2507 as the local LLM backbone in our framework. CBraMod is selected due to its SOTA performance among EEG foundation models and its pretraining on large, diverse EEG datasets. In addition, our encoder ablation study comparing CBraMod with LaBraM shows that CBraMod consistently achieves better performance. We chose Qwen-4B-Instruct-2507 as the LLM backbone because it outperforms other local LLMs in unimodal baselines and reliably follows instruction prompts to generate outputs with the required structure. In Table 5, we provide the hyperparameters of the alignment modules used in our study.
@@ C.3 Prompts
Figure 8 illustrates the prompt used in our approach
along with the EEG tokens. We also provide
examples of the sections in the prompt.
[…]
@@
[…]
@@
[…]
@@ D.2 Extended Report Generation Performance on S0002
Tables 8 and 9 present a comprehensive performance comparison of EEG report generation on the S0002 site under settings with and without patient clinical context. Both proposed variants consistently outperform all baselines across all evaluation metrics. Among them, CELM achieves better performance compared to the memory-efficient CELM -SCC variant across all metrics, except for BERTScore in the zero-context setting.
@@ D.4 EEG Encoder Ablation
We conducted an ablation study to analyze report generation performance across different EEG encoders. Specifically, we compared CBraMod (Wang et al., 2024b) and LaBraM (Jiang et al., 2024), both EEG foundation models pretrained on large-scale EEG datasets. The results show that CBraMod consistently outperforms LaBraM, highlighting the critical role of high-quality EEG representations in effective clinical report generation.
@@
D.5 Score Distribution
@@
D.6 More Results on Section-wise Analysis
Table 12: Performance by report section on S0001 dataset
Table 12 reports the complete section-wise performance across all models and evaluation metrics. These results extend the main text by providing detailed comparisons for each report section and further show the consistency of CELM across diverse clinical report components.
@@ D.7 Qualitative Analysis and Case Studies
This subsection presents qualitative analyses and representative case studies to complement the quantitative evaluation. In Figure 111213, and  14, we compare generated reports from the unimodal baseline, CELM-SCC, and CELM against ground-truth reports across multiple EEG report sections and datasets.

### 13 May 2026 v3

@@ Appendix C Implementation Details
To facilitate reproducibility, we provide the complete source code for constructing the EEG–report benchmark from the Harvard Electroencephalography Database and the EHR database hosted on the Brain Data Science Platform, together with our model implementation, training scripts, data loaders, and pretrained weights at https://anonymous.4open.science/r/CELM-3AF4. Additionally, we provide the detailed implementation details of our approach in this section and the prompts used for the baselines in Appendix B. The dataset utilized for this study is publicly available and can be accessed at https://bdsp.io/content/harvard-eeg-db/4.1/ after obtaining credentialed access. Experiments were conducted using multiple NVIDIA RTX 6000 and NVIDIA A100 GPUs with 49GB and 80GB of memory.
@@ C.1 Model Details and Hyperparameters
We use CBraMod [28] as the EEG encoder and Qwen-4B-Instruct-2507 as the local LLM backbone in our framework. CBraMod is selected due to its SOTA performance among EEG foundation models and its pretraining on large, diverse EEG datasets. In addition, our encoder ablation study comparing CBraMod with LaBraM shows that CBraMod consistently achieves better performance. We chose Qwen-4B-Instruct-2507 as the LLM backbone because it outperforms other local LLMs in unimodal baselines and reliably follows instruction prompts to generate outputs with the required structure. In Table 5, we provide the hyperparameters of the alignment modules used in our study.
@@ C.3 Prompts
Figure 10 illustrates the prompt used in
our approach along with the EEG tokens.
We also provide examples of the
sections in the prompt.
ELM Prompt Input : EEG projected tokens prepended to text tokens. EEG CHANNELS [’C3’, ’C4’, ’O1’, ’O2’, ’Cz’, ’F3’, ’F4’, ’F7’, ’F8’, ’Fz’, ’Fp1’, ’Fp2’, ’Fpz’, ’P3’, ’P4’, ’Pz’, ’T3’, ’T4’, ’T5’, ’T6’, ’A1’, ’A2’] You are an expert clinical neurophysiologist specializing in EEG interpretation and clinical report generation. TASK
@@
Patient history EEG description EEG Channels EEG SECTION DESCRIPTIONS [STANDARDIZED_SECTION_DESCRIPTIONS] e.g. EEG DESCRIPTION/DETAILS: Detailed narrative of EEG findings including background activity, sleep stages, physiologic variants, and abnormalities observed during the recording period. GUIDELINES Generate only the sections listed in **SECTIONS TO BE GENERATED**. Do NOT generate any additional sections. Do NOT repeat the same section more than once. Only generate the output in the JSON format and do not include any other text or explanation. OUTPUT FORMAT (STRICT)
@@
json {"report_sections": [ {"section_name": "Name of the section as given in SECTIONS TO BE GENERATED", "section_text": "Generated text for the section in string"}, ... ]} SECTIONS TO BE GENERATED
@@ D.2 Extended Report Generation Performance on S0002
Tables 8 and 9 present a comprehensive performance comparison of EEG report generation on the S0002 site under settings with and without patient clinical context (subset of S0002 samples without patient clinical context). Both proposed variants consistently outperform all baselines across all evaluation metrics. Among them, CELM achieves better performance compared to the memory-efficient CELM-SCC variant across all metrics, except for BERTScore in the zero-context setting.
@@ D.4 EEG Encoder Ablation
We conducted an ablation study to analyze report generation performance across different EEG encoders. Specifically, we compared CBraMod [28] and LaBraM [10], both EEG foundation models pretrained on large-scale EEG datasets. The results show that CBraMod consistently outperforms LaBraM, highlighting the critical role of high-quality EEG representations in effective clinical report generation.
@@
D.5 Performance Analysis by Report Section
Clinical EEG reports comprise multiple sections, each serving a distinct purpose: EEG description/details provides a detailed narrative of observed waveforms and patterns, impression/interpretation summarizes the clinical significance of findings, background activity characterizes baseline rhythms, and events/seizures documents seizure episodes. To evaluate whether CELM can reliably generate these diverse sections, we conduct a section-wise analysis on S0001, which exhibits richer section diversity compared to S0002 (where reports predominantly contain only EEG description/details; see Figure 7 in Appendix). Figure 11 presents ROUGE-1 scores, comparing CELM against the best-performing unimodal baseline (text + EEG features) from each LLM family. CELM achieves the highest performance in 6 out of 7 report sections. Performance degrades in the interictal epileptiform abnormalities section, highlighting a key limitation and a challenge for ELMs in modeling rare and clinically complex events, an important direction for future work. Detailed results are provided in Table 12, which reports the complete section-wise performance across all models and evaluation metrics. These results further demonstrate the consistency of CELM across diverse components of clinical reports.
Table 12: Performance by report section on S0001 dataset
D.6 Score Distribution
@@ D.7 Qualitative Analysis and Case Studies
Figure 14 presents qualitative comparisons on S0002 between the best-performing unimodal baseline (text with EEG features), the linear projector variant, and CELM, alongside ground truth reports. Figure 14a shows the EEG description/details section, while Figure 14b shows the impression/interpretation section. Across these examples, CELM produces reports that are more closely aligned with the ground truth, correctly identifying clinically relevant findings such as moderate bilateral beta activity, intermittent drowsiness, and bilateral slowing in the theta and delta ranges. In contrast, the linear projector and unimodal baselines fail to capture these. In the impression section, CELM identifies the recording as abnormal, whereas the linear projector variant incorrectly predicts it as normal. These results demonstrate the promise of ELMs in learning to generate clinical reports from unstructured notes paired with long EEG recordings. At the same time, they underscore the need for more rigorous and standardized benchmarking to assess these models. Additional examples spanning both sites, diverse report sections, and varying performance levels are provided below to further illustrate model strengths and failure modes. In Figure 151617, and  18, we compare generated reports from the unimodal baseline, CELM-SCC, and CELM against ground-truth reports across multiple EEG report sections and datasets.
@@
Appendix E Discussion, Limitations and Implications for Future work
Appendix F Broader Impact Statement